OpenStudy (anonymous):

Find an equation for the function f that has the given derivative and whose graph passes through the given point. Derivative Point f '(x) = 3 sin(9x) ( π/9 , − 14/3) can someone please help me?

OpenStudy (anonymous):

derivitive...f'(x)=3sin(9x) Point....(pi/9,-14/3)

OpenStudy (anonymous):

Methinks you're looking for something like this: \[\int\limits3\times \sin(9x)dx\] And then you have to tweak the constant so that it passes through the given point.

OpenStudy (anonymous):

hmmm oki im still conffused could you work me through it?:S

OpenStudy (anonymous):

i won't work you through the integration, i (my calculator) states that \[\int 3 \times sin(9x)dx = -\frac{1}{3}cos(9x)+C\] Now \[f(\frac{\pi}{9})=\frac{-14}{3}\] \[-\frac{1}{3}cos(9\frac{\pi}{9})+C=\frac{-14}{3}\] You should be able to continue from here on.

OpenStudy (anonymous):

The solution is: \[f(x)=-\frac{1}{3}cos(9x)-5\]

OpenStudy (anonymous):

thank you so much this actually amkes sense..thanks...is it minus 5 its hard to read it

OpenStudy (anonymous):

yes, -5, minus 5